OpenStudy (anonymous):

multiply (2 sqrt 3 - 1)(3 sqrt 3 - 2)

OpenStudy (kuoministers):

use FOIL ever heard of it?

OpenStudy (kuoministers):

Here is an example of FOIL (4x + 6)(x + 2) First - multiply the first term in each set of parenthesis: 4x * x = 4x2 Outside - multiply the two terms on the outside: 4x * 2 = 8x Inside - multiply both of the inside terms: 6 * x = 6x Last - multiply the last term in each set of parenthesis: 6 * 2 = 12

OpenStudy (anonymous):

yes

OpenStudy (kuoministers):

\[(2\sqrt{3} - 1)(3\sqrt{3} - 2)\]

OpenStudy (kuoministers):

can you complete it now?